OpenStudy (yato_gami):

I need help with this please! I want someone to go through the steps with me, not tell me the answer. Thanks~! ^_^ The price of bronze has increased by 10% per year from 2000. In the year 2000, Harry bought a bronze medal for $120. Which of the following functions f(x) can be used to represent the price of the medal x years after 2000?

OpenStudy (yato_gami):

A. f(x) = 120(0.9)x B. f(x) = 10(120)x C. f(x) = 120(1.1)x D. f(x) = 10(1.1)x

OpenStudy (sunnnystrong):

Okay so: Recall that \[A=A _{0}(b)^t\] A= amount Ao= initial b= growth factor (1+ percent of growth) t= time in years

OpenStudy (sunnnystrong):

Any ideas @Yato_Gami ?

OpenStudy (yato_gami):

Um I'm not too sure. :(

OpenStudy (sunnnystrong):

So it looks like Harry bought a piece of bronze medal in 2000 for $120. aka Ao=$120 Since then... the price has started to increase by 10% each year after 2000 If b= growth factor (1+ percent of growth) & 10% is the percent of growth b=1.10 So: What is the equation?

OpenStudy (yato_gami):

A = 120(1.1)t?

OpenStudy (sunnnystrong):

Yep :D looks good!

OpenStudy (yato_gami):

Oh ok! Is that it?